# Tropical Gradient Descent

### v1

@@ 4 Numerical Experiments
In this section we compare practical performance of classical descent, tropical descent, SGD, tropical SGD, Adam, Adamax, and Trop Adamax, with a focus on their stability behaviour. The implementation for all these experiments is available in an anonymous GitHub repository at https://github.com/Roroast/TropicalGradDescent.
@@ 4.1 Methods
Descent
We will perform classical descent (CD) and tropical descent (TD) using equivalent same step size sequences $a_{n}$ for both methods, which we fix as
@@
Stochastic Gradient Descent
Stochastic gradient descent is a variant of gradient descent which looks to improve computational speed by approximating $\nabla f(t)$; rather than computing the loss and its gradient at $\mathbf{t}$ with respect to the entire dataset $f(\mathbf{t};X)$, we evaluate the loss and its gradient with respect to a random sample $\mathbf{x}_{n}\in X$, $f(\mathbf{t};\mathbf{x}_{n})$. This is simplified with the notation $f_{n}(\mathbf{t})$. As with gradient descent, we use the step size sequence give by
@@
Adam Variants
The Adam algorithm (Kingma & Ba, (2017)) uses gradient momentum to compound consistent but small effects. This is done by taking first and second moments of the gradient over all steps using exponentially decaying weights. The Adamax algorithm is a variant of the Adam algorithm motivated by the use of the $\infty$-moments rather than second moments.
To adapt the Adam algorithm for the tropical setting, we take moments of an un-normalized tropical descent direction. When looking to take higher moments of $\mathbf{m}_{n}$ over $n$, the $L_{\infty}$ is the natural option for the tropical setting. We therefore look to use the Adamax recursion relation on the $\infty$-moment estimate $\mathbf{u}_{n}$:
$\mathbf{u}_{n+1}=\max(\beta_{2}\mathbf{u}_{n},\mathbf{d}_{n}).$
We therefore define TropAdamax as in Algorithm 2.
Algorithm 2 TropAdamax
@@
     $\mathbf{g}_{n}\leftarrow\nabla f(\mathbf{t}_{n-1})$;
     $\mathbf{d}_{n}\leftarrow[\mathbf{g}_{n}>0]*(\max(\mathbf{g}_{n})-\min(\mathbf{g}_{n}))$;
@@ 4.2 Computation
We test the gradient methods discussed on 12 datasets of different sizes, shapes and dimensions; we consider datasets generated by a branching process, a coalescent process, or a Gaussian distribution on $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, for $N=6,28$ and samples of size 10 or 100. See Appendix A for further details on the simulated datasets. The learning rates used have been tuned for each gradient method and dataset size and dimension (Appendix C). In all experiments, random initialisations are taken from a Gaussian distribution on $\mathbb{R}^{N}$. We provide an overview of the general behaviour of each method for the various optimisation problems introduced in section 2.3.
Rather than investigating the relative convergence rates of each method, we are interested in the relative stability of local minima. We therefore plot the CDF of the relative log error after 1000 steps over 50 random initialisations, which can highlight certain error values at which several initialisations have accumulated. When the CDF of one method dominates another at log error $=\epsilon$, we expect it has a higher probability of achieving a log error of at most $\epsilon$.
@@ 4.2.1 Centrality Statistics
We first consider the behavior of tropical descent methods for the centrality statistics, Fermat–Weber points and Fréchet means. These optimisation problems are classically convex problems, as well as being both $\max$-tropical and $\min$-tropical location problems, and therefore lend themselves well to gradient methods. The experimental behaviour of these problems is similar, so in this section we include the results for the Fermat–Weber problem while the results for Fréchet means can be found in Appendix E.
Figure 5 shows the CDF of the final log relative error of the non-stochastic methods for the 12 datasets. We note that the stochastic methods achieves the worst error for all 12 datasets, and TropSGD is often performing worse than SGD. The disparity between SGD and TropSGD is greatest when the dimensionality of the data is greater than size of the dataset, while the disparity is smallest when the dimensionality is much smaller than the dataset size. We see similar behaviour in the log error of Tropical Descent and TropAdamax; they perform on par with their classical counterparts for each dataset other than the 28 by 10 datasets, but when the data dimensionality exceeds the number of data points, the tropical gradient methods are not achieving the same accuracy. We also note that convergence of all methods is less consistent for coalescent data than branching or Gaussian data; while it is unclear why this is, we note that coalescent data is supported on the space of ultrametric trees, which is a tropical hyperplane (Ardila & Klivans, (2006); Page et al., (2020)) and may influence the uniqueness of minima.
Figure 5: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the Fermat–Weber objective. Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
Data CD TD SGD TropSGD Adam Adamax TropAdamax
6$\times$10 Branching Data -4.60 -4.60 -4.48 -2.95 -4.60 -4.60 -4.60
6$\times$10 Coalescent Data -4.58 -4.55 -4.38 -1.34 -4.58 -4.58 -4.55
6$\times$10 Gaussian Data -4.60 -4.60 -4.44 -2.63 -4.58 -4.58 -4.60
6$\times$100 Branching Data -4.60 -4.60 -4.45 -3.95 -4.59 -4.59 -4.59
6$\times$100 Coalescent Data -4.55 -4.52 -4.37 -2.76 -4.54 -4.57 -4.54
6$\times$100 Gaussian Data -4.60 -4.60 -4.45 -4.00 -4.59 -4.60 -4.59
28$\times$10 Branching Data -4.60 -4.60 -4.15 -1.90 -4.60 -4.60 -4.56
28$\times$10 Coalescent Data -4.52 -3.76 -3.99 -2.24 -4.52 -4.56 -3.84
28$\times$10 Gaussian Data -4.59 -4.59 -4.09 -2.47 -4.60 -4.60 -4.47
28$\times$100 Branching Data -4.59 -4.59 -4.05 -3.13 -4.59 -4.59 -4.59
28$\times$100 Coalescent Data -4.49 -4.47 -3.86 -2.54 -4.38 -4.56 -4.48
28$\times$100 Gaussian Data -4.59 -4.59 -4.09 -3.02 -4.59 -4.59 -4.56
Table 1: The mean log relative error of each gradient method after 1000 steps for each dataset when minimising the Fermat–Weber objective function. The minimal mean errors for each dataset are in bold.
@@ 4.2.2 Wasserstein Projections
The Wasserstein projection problems highlight the differences in convergent behaviour with and without the tropical gradient, as they are $\max$-tropical location problems, but not classically convex. The CDFs of the final log relative errors can be found in Figures 6 and 7.
Figure 6: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the 2-Wasserstein projection objective. Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
Figure 7: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the $\infty$-Wasserstein projection objective. Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
In Figures 6 and 7 we see large steps at various intervals in the CDF plots, which is indicative of stable local minima. This is particularly prevalent in low dimensional space. Classical descent, Adam, and Adamax are particularly susceptible to this, while tropical descent and TropAdamax are often more likely to pass these local minima. This difference is most pronounced when solving the $\infty$-Wasserstein problem in low dimensions—we are almost certain to reach the minimum with small error, while classical methods have a 50-80% chance of achieving the same error.
Data CD TD SGD TropSGD Adam Adamax TropAdamax
6$\times$10 Branching Data -3.52 -4.13 -1.98 -2.17 -3.49 -3.51 -3.99
6$\times$10 Coalescent Data -3.53 -4.31 -2.38 -3.17 -3.77 -3.85 -4.16
6$\times$10 Gaussian Data -3.66 -3.72 -3.21 -2.76 -3.84 -3.85 -3.90
6$\times$100 Branching Data -3.57 -4.38 -3.28 -1.64 -3.94 -3.94 -4.39
6$\times$100 Coalescent Data -4.35 -4.54 -3.31 -2.54 -4.43 -4.44 -4.26
6$\times$100 Gaussian Data -3.78 -4.15 -3.39 -2.69 -3.77 -3.79 -4.42
28$\times$10 Branching Data -3.59 -4.29 -3.18 -2.03 -3.63 -3.64 -4.34
28$\times$10 Coalescent Data -4.07 -3.80 -3.45 -2.50 -4.12 -4.09 -3.82
28$\times$10 Gaussian Data -3.53 -3.89 -3.22 -2.18 -3.61 -3.58 -3.85
28$\times$100 Branching Data -3.65 -4.54 -3.42 -2.64 -3.86 -3.87 -4.53
28$\times$100 Coalescent Data -4.12 -4.57 -3.89 -3.44 -4.20 -4.21 -4.55
28$\times$100 Gaussian Data -3.65 -4.50 -3.39 -2.67 -3.85 -3.86 -4.48
Table 2: The mean log relative error of each gradient method after 1000 steps for each dataset when minimising the $2$-Wasserstein projection objective function. The minimal mean errors for each dataset are in bold.
Data CD TD SGD TropSGD Adam Adamax TropAdamax
6$\times$10 Branching Data -2.56 -4.29 -0.58 -0.94 -2.58 -2.59 -4.28
6$\times$10 Coalescent Data -4.07 -4.48 -1.13 -1.38 -4.13 -4.09 -4.48
6$\times$10 Gaussian Data -4.00 -4.44 -2.08 -3.10 -4.04 -4.02 -4.38
6$\times$100 Branching Data -3.63 -4.46 -2.14 -1.17 -3.60 -3.65 -4.46
6$\times$100 Coalescent Data -4.11 -4.51 -2.19 -2.30 -4.18 -4.14 -4.51
6$\times$100 Gaussian Data -3.02 -4.36 -1.48 -1.23 -3.06 -3.09 -4.31
28$\times$10 Branching Data -3.72 -3.66 -2.19 -1.14 -3.81 -3.81 -3.55
28$\times$10 Coalescent Data -4.31 -4.13 -2.83 -1.99 -4.39 -4.39 -4.01
28$\times$10 Gaussian Data -4.16 -3.47 -2.08 -0.96 -4.43 -4.36 -3.12
28$\times$100 Branching Data -3.55 -3.86 -1.38 -1.16 -3.59 -3.49 -3.71
28$\times$100 Coalescent Data -4.45 -4.12 -2.11 -2.05 -4.54 -4.54 -3.95
28$\times$100 Gaussian Data -3.23 -3.61 -1.74 -1.61 -3.27 -3.22 -2.94
Table 3: The mean log relative error of each gradient method after 1000 steps for each dataset when minimising the $\infty$-Wasserstein projection objective function. The minimal mean errors for each dataset are in bold.
From tables 2 and 3, we see that when solving the 2-Wasserstein projection problem, tropical methods are optimal in all cases other than $28\times 10$ coalescent data. For the $\infty$-Wasserstein projection problem in small dimensions, tropical descent outperforms other methods by some margin, while the comparative performances are less consistent for the $\infty$-Wasserstein projection problem in high dimensions.
@@ 4.2.3 Linear Regression
The linear regression problem is a $\min$ tropical location problem, so we run the same code with input $-\mathbf{t}$. As the coalescent data lies on a tropical hyperplane (Ardila & Klivans, (2006); Page et al., (2020)), the minimum of the linear regression function for such data is exactly 0. Figure 8 therefore shows the CDF of the log error rather than the CDF of the log relative error.
@@
As we can see in table 4, the tropical methods outperform their classical counterparts for every dataset, often by a factor of two. The tropical descent method in particular is the most effective for all but one dataset. In Figure 8 we see that the CDF of log errors for tropical descent and TropAdamax consistently dominates those of classical descent, Adam and Adamax, to the extent that the support of their CDFs does not intersect that of their classical counterparts for some datasets. We also note that unlike the previous problems, TropSGD is outperforming SGD when solving the linear regression problem, often to the extent that TropSGD outperforms all classical gradient methods.
@@
Figure 8: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the linear regression objective. Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
Data CD TD SGD TropSGD Adam Adamax TropAdamax
6$\times$10 Branching Data -0.96 -5.96 -0.92 -2.10 -1.09 -1.07 -5.57
6$\times$10 Coalescent Data -4.27 -6.25 -3.80 -5.78 -4.44 -4.54 -6.22
6$\times$10 Gaussian Data -1.13 -5.42 -0.86 -1.33 -1.15 -1.14 -4.79
6$\times$100 Branching Data -1.08 -5.34 -0.86 -1.54 -1.10 -1.11 -3.73
6$\times$100 Coalescent Data -6.42 -6.35 -3.29 -5.69 -6.81 -6.88 -6.22
6$\times$100 Gaussian Data -0.93 -5.43 -1.09 -1.93 -0.92 -0.92 -3.97
28$\times$10 Branching Data -1.60 -5.83 -1.13 -4.47 -1.41 -1.48 -5.53
28$\times$10 Coalescent Data -2.56 -6.52 -2.22 -5.50 -2.68 -2.61 -6.44
28$\times$10 Gaussian Data -1.35 -5.68 -0.91 -3.99 -1.20 -1.23 -5.56
28$\times$100 Branching Data -0.79 -4.33 -0.54 -1.46 -0.69 -0.70 -3.44
28$\times$100 Coalescent Data -3.35 -5.34 -2.10 -6.04 -3.79 -3.94 -4.95
28$\times$100 Gaussian Data -0.93 -3.88 -0.64 -1.22 -0.84 -0.85 -3.44
@@
Table 4: The mean log relative error of each gradient method after 1000 steps for each dataset when minimising the linear regression objective function. The minimal mean errors for each dataset are in bold.
5 Discussion

### v2

@@ 4 Numerical Experiments
In this section we compare the practical performance of classical descent, tropical descent, SGD, tropical SGD, Adam, Adamax, and TrAdamax, with a focus on their stability behavior. The implementation for all these experiments is available in the GitHub repository at https://github.com/Roroast/TropicalGradDescent.
@@ 4.1 Methods
We begin by outlining our implementation of steepest descent, stochastic gradient descent (SGD), and Adam as well as their tropical variants. We take a heuristic approach to the tropicalization of SGD and Adam, replacing derivatives with tropical steepest descent directions in each case.
Steepest Descent. We will perform classical descent (CD) and tropical descent (TD) using equivalent step size sequences $a_{n}$ for both methods, which we fix as
@@
Stochastic Gradient Descent. Stochastic gradient descent is a variant of gradient descent which looks to improve computational speed by approximating $\nabla f(t)$; rather than computing the loss and its derivative at $\mathbf{t}$ with respect to the entire dataset $f(\mathbf{t};X)$, we evaluate the loss and its derivative with respect to a random sample $\mathbf{x}_{n}\in X$, $f(\mathbf{t};\mathbf{x}_{n})$. This is simplified with the notation $f_{n}(\mathbf{t})$. As with gradient descent, we use the step size sequence give by
@@
Adam Variants. The Adam algorithm (Kingma & Ba,, 2015) uses gradient momentum to compound consistent but small effects. This is done by taking first and second moments of the gradient over all steps using exponentially decaying weights. The Adamax algorithm is a variant of the Adam algorithm motivated by the use of the $\infty$-moments rather than second moments.
Algorithm 2 TrAdamax
@@
     $\mathbf{g}_{ni}\leftarrow\nabla f_{i}(\mathbf{t}_{n-1})$;
     $\mathbf{d}_{ni}\leftarrow(\max_{j}(\mathbf{g}_{nj})-\min_{j}(\mathbf{g}_{nj}))*\mathbf{1}_{\mathbf{g}_{ni}>0}$;
@@
To adapt the Adam algorithm for the tropical setting, we take moments of an unnormalized tropical descent direction. When looking to take higher moments of the gradient over $n$, the $L_{\infty}$ is the natural option for the tropical setting. We therefore look to use the Adamax recursion relation on the $\infty$-moment estimate $\mathbf{u}_{n}$:
$\mathbf{u}_{n+1}=\max(\beta_{2}\mathbf{u}_{n},\mathbf{d}_{n}).$
The tropicalized Adamax algorithm is then given by Algorithm 2.
@@ 4.2 Computation
We now look to solve our centrality statistics, Wasserstein projection problems, and tropical linear regression using the 7 gradient methods discussed: classical descent (CD), tropical descent (TD), stochastic gradient descent (SGD), tropical stochastic gradient descent(TSGD), Adam, Adamax, and TrAdamax. We run computations for 12 datasets of different sizes, shapes and dimensions; motivated by phylogenetic data science applications, we consider datasets generated by branching processes, a coalescent processes, or Gaussian distributions on $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, for $N=6,28$ and samples of size 10 or 100. We take same random initializations for each gradient method, taken from a Gaussian distribution on $\mathbb{R}^{N}$. Appendix B provides further details on the simulated datasets and computational methodology used. For each gradient method, we use learning rates which have been optimized for each dataset size and dimension, across different distribution types (Appendix C).
Rather than investigating the relative convergence rates of each method, we are interested in the relative stability of local minima. We therefore plot the CDF of the relative log error after 1000 steps over 50 random initializations, which can highlight certain error values at which several initializations have accumulated. When the CDF of one method dominates another, it has a higher probability of achieving the same error and is therefore a more suitable method.
As we will observe, tropical gradient methods are stable at fewer points than classical gradient methods; as a result, tropical methods typically have further to travel before reaching a stable point. Over few steps, this would be compensated for by selecting larger learning rates for tropical gradient methods relative to classical gradient methods, leading to less precise results in the case of tropical gradient methods. Our computations are run over 1000 steps, as this gives enough time for tropical gradient methods to stabilise, given our initialisations, without sacrificing learning rate; we see in Appendix C that neither classical nor tropical learning rates are consistently greater than the other.
@@ 4.2.1 Centrality Statistics
We first consider the behavior of tropical descent methods for centrality statistics, Fermat–Weber points and Fréchet means. These optimization problems are classically convex problems, as well as being both $\max$-tropical and $\min$-tropical location problems, and therefore lend themselves well to gradient methods. The experimental behavior of these problems is similar, so in this section we include the results for the Fermat–Weber problem while the results for Fréchet means can be found in Appendix E.
Figure 5 shows the CDF of the log relative error over 1000 steps of each gradient method for the 12 datasets. We note that the deterministic methods achieve similar results for most datasets, while the stochastic methods achieve the worst error for all 12 datasets and TSGD is performing worse than SGD. The disparity between SGD and TSGD is greatest when the dimensionality of the data is greater than the size of the dataset, and for coalescent data. We see similar behavior in the log error of Tropical Descent and TrAdamax; they perform on par with their classical counterparts for each dataset other than the 28 by 10 coalescent data, for which the tropical gradient methods are not achieving the same accuracy. While it is unclear why this is, we note that coalescent data is supported on the space of ultrametric trees, which is a tropical hyperplane (Ardila & Klivans,, 2006; Page et al.,, 2020) and may influence the uniqueness of minima.
Figure 5: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the Fermat–Weber objective. Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
Table 1: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the Fermat–Weber objective function. The minimal mean errors for each dataset are in bold.
Data CD TD SGD TSGD Adam Adamax TrAdamax
6$\times$10 Branching Data -4.60 -4.60 -4.50 -3.32 -4.60 -4.60 -4.60
6$\times$10 Coalescent Data -4.58 -4.55 -4.40 -1.28 -4.57 -4.57 -4.57
6$\times$10 Gaussian Data -4.60 -4.60 -4.44 -2.75 -4.60 -4.60 -4.56
6$\times$100 Branching Data -4.60 -4.60 -4.39 -4.02 -4.54 -4.58 -4.59
6$\times$100 Coalescent Data -4.55 -4.54 -4.25 -2.88 -4.56 -4.58 -4.55
6$\times$100 Gaussian Data -4.60 -4.60 -4.39 -4.03 -4.52 -4.57 -4.59
28$\times$10 Branching Data -4.60 -4.60 -4.15 -1.89 -4.60 -4.60 -4.56
28$\times$10 Coalescent Data -4.53 -3.76 -3.96 -2.11 -4.52 -4.57 -3.87
28$\times$10 Gaussian Data -4.58 -4.58 -4.07 -2.40 -4.60 -4.60 -4.47
28$\times$100 Branching Data -4.59 -4.57 -4.04 -2.89 -4.59 -4.59 -4.59
28$\times$100 Coalescent Data -4.51 -4.54 -3.91 -2.33 -4.39 -4.58 -4.39
28$\times$100 Gaussian Data -4.59 -4.53 -4.10 -2.78 -4.59 -4.59 -4.59
@@ 4.2.2 Wasserstein Projections
The Wasserstein projection problems highlight the differences in convergent behavior between classical and tropical gradient methods, as they are $\min$-tropical location problems but not classically convex. The CDFs of the final log relative errors can be found in Figures 6 and 7.
Figure 6: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the 2-Wasserstein projection objective. Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
Figure 7: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the $\infty$-Wasserstein projection objective. Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
In Figures 6 and 7 we see large steps at various intervals in the CDF plots, which is indicative of stable local minima. This is particularly prevalent in low dimensional space. Classical descent, Adam, and Adamax are particularly susceptible to this, while tropical descent and TrAdamax are more likely to pass these local minima. This difference is most pronounced when solving the $\infty$-Wasserstein problem in low dimensions—we are almost certain to reach the minimum with small error, while classical methods have a 50-80% chance of achieving the same error.
@@ 4.2.3 Linear Regression
As the coalescent data lies on a tropical hyperplane (Ardila & Klivans,, 2006; Page et al.,, 2020), the minimum of the linear regression function for such data is exactly 0. Figure 8 therefore shows the CDF of the log error rather than the CDF of the log relative error.
@@
Table 2 shows the mean log errors of each gradient method and dataset, from which we see that the tropical methods outperform their classical counterparts for all but one dataset, often by a factor of two. In Figure 8 we see that the CDF of log errors for tropical descent and TrAdamax consistently dominates those of classical descent, Adam and Adamax, to the extent that the support of their CDFs does not intersect that of their classical counterparts for some datasets. We also note that unlike the previous problems, TSGD is outperforming SGD when solving the linear regression problem, often to the extent that TSGD outperforms all classical gradient methods.
@@
Figure 8: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the linear regression objective. Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
@@
Table 2: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the linear regression objective function. The minimal mean errors for each dataset are in bold.
Data CD TD SGD TSGD Adam Adamax TrAdamax
6$\times$10 Branching Data -0.97 -5.79 -0.56 -2.16 -0.80 -0.84 -4.39
6$\times$10 Coalescent Data -2.70 -5.87 -2.11 -5.47 -2.97 -2.66 -6.58
6$\times$10 Gaussian Data -1.19 -5.48 -0.80 -1.35 -1.05 -1.10 -4.10
6$\times$100 Branching Data -0.79 -5.49 -0.71 -1.43 -0.72 -0.72 -3.53
6$\times$100 Coalescent Data -5.54 -5.76 -3.11 -5.57 -6.90 -7.01 -6.52
6$\times$100 Gaussian Data -1.04 -5.50 -0.88 -1.70 -1.02 -1.02 -3.81
28$\times$10 Branching Data -1.49 -5.51 -0.93 -4.49 -1.14 -1.14 -5.57
28$\times$10 Coalescent Data -2.43 -6.63 -1.77 -5.55 -2.16 -2.29 -6.18
28$\times$10 Gaussian Data -1.26 -5.16 -0.79 -4.29 -1.08 -1.08 -5.72
28$\times$100 Branching Data -0.90 -4.37 -0.61 -1.42 -0.85 -0.77 -3.32
28$\times$100 Coalescent Data -3.33 -5.84 -2.17 -5.35 -3.02 -3.20 -4.93
28$\times$100 Gaussian Data -0.91 -4.35 -0.62 -1.07 -0.96 -0.90 -3.51
5 Conclusion and Discussion
We have introduced the concept of tropical descent as the steepest descent method with respect to the tropical norm. We have demonstrated, both theoretically and experimentally, that it possesses well-behaved properties with respect to a wide class of optimization problems on the tropical projective torus, including several key statistical problems for the analysis of phylogenetic data. As well as proposing our own framework for tropical gradient methods, this work also provides a detailed review of the varying behavior of popular gradient methods when applied to tropical statistical optimization problems.